# Identifying and Analyzing Performance-Critical Tokens in Large Language Models

### v3

@@ Appendix V Supplemental Experiments for the Structural Cue Characteristic
During LLM pre-training, it is very likely that the model has seen text formatted in a similar way (e.g. in html, plain text with headings), in which the LLM would learn to recognize and store information in the representation of these formatting tokens. A recent study also provides supporting facts from the perspective of pretraining for this (Chen et al., 2024).
@@
<m>: [delimiter: random template 1] |
<m><m><m> … <m> [masked demonstration] |
<m>: [delimiter: random template 2] |
@@
Datasets | Notations | Examples |
Randomfixed | ||
@@
Datasets | Notations | Examples |
Randomfixed | ||
@@
Datasets | Notations | Examples |
Randomfixed | ||
@@
Datasets | Notations | Examples |
Randomfixed | ||
@@
Datasets | Notations | Examples |
Randomfixed | ||

### v4

@@ Appendix X Discussion about the Characteristic of Structural Cue
During LLM pre-training, it is very likely that the model has seen text formatted in a similar way (e.g. in html, plain text with headings), in which the LLM would learn to recognize and store information in the representation of these formatting tokens. A recent study also provides supporting facts from the perspective of pretraining for this (chen2024parallel).
@@
| (m): [delimiter: random template 1] |
| (m)(m)(m) … (m) [masked demonstration] |
| (m): [delimiter: random template 2] |
@@
| Datasets | Notations | Examples | |
| Randomfixed | |||
@@
| Datasets | Notations | Examples | |
| Randomfixed | |||
@@
| Datasets | Notations | Examples | |
| Randomfixed | |||
@@
| Datasets | Notations | Examples | |
| Randomfixed | |||
@@
| Datasets | Notations | Examples | |
| Randomfixed | |||